UT 14 Sep 2021




Night Staff:

OSA: D.G. Huerta

ISp: C. Bolyard ISp: ISA/Observers: J.Power

Technical Wizard: J. Hill (Tucson)

Mtn Manager: M. Wagner

Sunset: 18:33

Twlight Ends: 19:51

Twilght Starts: 04:40

Sunrise: 05:58

Starting Config: LUCI-Bino (LUCI-1 authorized)

Night Log:


Proposed Plan of Action:

  • LUCI-1:
    • DONE LUCI-1 Dual BLR target - twilight until 30deg limit (Rothberg) - until 20:45 local time
      • Take Telluric first as soon as AGW-1 can wave-front sense
      • If opening is delayed, until 20hr - configure for MODS and take observations with MODS (if available)
  • Reconfigure for MODS-Bino - skip if we have transform ready to implement
    • Keep transform in place
    • Rho Oph - observations
    • Spectrophotometric Standard
  • If Transforms ready proceed to appropriate instrument to collect AGW-1 Transform Confirmatory data - highest prioriy
  • LUCI1:
    • DONE AGW-1/2 Transform Confirmatory Data - 1.5hr - Highest Priority
      • Fields:
        • StoneO 21 37 31 +00 31 55 Stone_O_m12.list - airmass 1.3 at 20:45
        • StoneP 22 35 49 +01 31 33 Stone_P1.list
        • StoneA 00 54 02 -00 32 34 Stone_A3.list
      • Take Confirmatory Rot ZP through Astrometric masks (note: the Rot ZP was not updated for LUCI-1 it was determined to be OK as-is)
    • Photometric Standards (clear any seeing)
    • DONE NS dither test - fast target
  • MODS - reconfigure after NS longslit test unless seeing is so poor - then move to LBCs
    • Dither Test - M52 or M71 - high priority
* Spectrophotometric stars (if clear any seeing) - all modes * Spectrophotometric stars (if clear any seeing) * Standard Stars (if clear any seeing)
    • Pointing Model (poor weather and poor seeing)
    • NS Longslit - low priority
    • Photometric and Spectro-photometric standard (if clear and poor seeing)
    • Spectrophotometric standard all modes
    • Grating: Dual, Blue, Red
    • Prism: Dual, Blue, Red
  • Mixed Mode Test - if time
    • LBC-Red & MODS-1
      • Test whether LBC can send a 2nd script (Dec -90) and everything works well for both sides
  • LBC:
    • Pointing Model LBC (poor seeing, clouds) - 2-3hrs - highest priority for LBCs - 02hr start pointing model latest
      • Turn on PCS logging
      • Turn on TMS passive mode
      • Need to collimate at each star - or - collimate every few stars
* Chip 2 only exposure * Chip 2 only exposure ( adjusted to be full frame) * File Location:
        • /home/lbto/Calib_OBs/PtStars_10m_OBs/VV
        • PtStarrs: /home/lbto/POINTING/PtStars_10m/PtStars_10m.dat
    • Photometric standards (any seeing - clear)
    • Spectrophotometric standards (any seeing - clear)
    • If conditions too poor for other stuff - Pointing Model in Active mode too
    • General TMS Active mode testing (any conditions)
      • run any Science-like OB in the LBC Restart OT file
      • Run for 1-1.5 hours
  • Poor Weather (if pointing models are exhausted)
    • LUCI-1 local mergers
    • Service_Observing/Restart_2021/LUCI/PoorWeather/COBandhead_Mergers_V17.xml - OT file
    • finder charts in sub-directory (but they are so large you can't miss 'em)

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - issues with M1RC headers & filenumbers short 1 digit
      • AGW-5 - available
    • MODS-2 - offline - glycol leak - to be inspected today
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - available
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
        • 8k - COMPLETED
      • 1x2
      • 2x2
    • Grating
      • Dual
        • 0.8" slit, 0.3" slit, slitless pixel flats, arcs - COMPLETED
      • Red
        • 0.8" slit, slitless pixel flats, arcs - COMPLETED
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
      • Narrow, medium-band filters
      • Broad zJHKKs and HeI completed
    • N1.8 G210 CWL=2.3 K-band 0.5" cals
    • N1.8 G200 CWL=1.93 HKspec 0.5" cals
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

Night Overview

1:20 Telescope has been released. David is reconfiguring for binocular luci, authorized LUCI1 only. Olga has checked things out post MODS2 glycol leak.
"The MODS2 test flats look 'clean as a whistle' (cleaner than I have ever seen them). I will dig up some flats for ratios - meanwhile I will take a few more test images - this time repeating a set of imaging cals I did a few days ago." She is taking some mods2 red only r images now while David reconfigures.

David has set the 26 el limit and has restarted GCS for the MODS wfs image

1:40 David is opening

1:48 Sending preset for Sky flats. Might be a bit late but we will see. We missed the window with the slightly late opening. Will try in the morning or tomorrow.

Dual BRL

1:55 David is moving for pointing and collimation near the tellurci for Dual BRL

1:59 GCS right just randomly went down while David was executing the pointing check.

2:14 Sending Preset to HIP 71172. Glenn is online to watch for LUCI bugs.

Seeing 0.65"-0.75". X-Shift -9.8328 Y-shift -0.4783

luci1.20210914.0018-19

2:26 Slewing to Dual BRL Target.

Offsets calculated -0.4155,-0.2839,0.4612, additional X of 0.14 required.

Seeing is 11.35" on the guider, 1.43" on the DIMM

Starting science luci1.20210914.0030

Seeing highly variable bouncing between 1.1-2.0". Determined the seeing is not stable enough sor good enough for transorm at this time so will be proceeding with the LUCI Stress test after this.

LUCI Stress Test Software

3:54 Slewing to Neptune for LUCI Stress test (NS preset to 899)

4:08 Saving Neptune image for a very long time. 45 seconds or so to save?
email alert: New data file luci1.20210914.0051.fits has reached a wait eof timeout.
file with 145 cube is 2.3 Gb so too big.

4:27 Sending same Neptune script but with 1x145x0.0 sec integrated. 17 MB

4:57 Sending subwindowed 256x256 1x145x0.31sec cube, 37M size images. Rtd running smoothly

5:09 Testing the subwindow tab in the rtd. Looks like the send for subwindow is not wired.

5:15 Manually set up for center stripe. Seeing is improving. 1". Will take some more images to see if it remains stable and also to demonstrate RtD slowdowns with many images. 291M image size

Aladin is not frozen but RTD is frozen. Glenn is seeing all the fun bugs. Aladin became completely non responsive on image 70. The aladin window vapour locked slowly.

LUCI Tranform Verification

5:33 Authorizing binocularly and sendning preset to Stone P for pointing and collimation

my bad, MONOCULARLY. No need for binocular for transform!
  • AGW Transform Confirmatory Data - 1.5hr - Highest Priority
    StoneP 22 35 49 +01 31 33 Stone_P1.list
  1. Updated oacontrol.conf with numbers provided but doug
  2. rdwrconfig -u1
  3. restart GCS
  4. configured luci with N30 field stop, H, flexure on.
  5. Start field_collect, EXP=4.0
first image luci1.20210914.0074

Somethign is not right. Not progressing, hung asking for a go on each exposure. David is cancelling the preset and we are starting again.

luci1.20210914.0076-0097

6:16 Seeing has dramatically improved. 0.65"-0.85"

6:37 Seeing start to oscillate a little 0.9-1.1"

NS Longstit LUCI Test

6:49 Slewing to 2021 LH9 for fast moving NS target longslit test. Target is pretty faint. This may not be the best for this test. Taking some test images but may abandon if spectrum not obvious.

6:13 Science started

Its there clearly in pair subtraction, monitoring. Other stars are flying through

luci1.20210914.014-0120

8:00 Slewing to zenith to reconfigure for bino LBC's

MISSED THE ROT ZERO POINT VERIFICATION THROUGH ASTROMETRIC MASK

LBC Pointing Model

Andrew is turning on the lasers

08:22 Starting Pointing log, sending first preset to a higher target for start of night collimation. First point not actually part of the series but it is always easier to collimate the first time ~70 degrees. First preset got copointing error. David restarted PCS. Restarted PCS.

Send but there seems to be a problem with the script :

4 2021/09/14 08:38:53.774981 OB execution error: unknown filter required [src/lbc/actions.c:1162]
3 2021/09/14 08:38:44.624034 >>> preset/expose/save interrupted at preset time, retcode=10
2 2021/09/14 08:38:37.939945 focus position for filter "V-BESSEL " not found on focus positions database "/lbccontrol/current/conf/leftchannelfocus.dat" [src/lbc/actions.c:1474]
1 2021/09/14 08:38:32.125768 filters system wrong preset request
0 2021/09/14 08:38:30.811626 FILTERS WHEEL#2 empty position or required filter not found >V-BESSEL < [src/filters/filters.c:14

tried a different filter but the problem persisted. Looks like we are doing this the old fashioned way. Remade scripts from scratch with OT

[_found later on that the problem was due to an extra space after the filter name. Removed that. opk 6/2/2022_]

Order of operations:

point at Pt Stars recommentded target, collimate, co-point with 1sec, run 20 sec sci script for field abberation analysis later, move on. TMS runnning passively. Sert point set after copointing on Star #8.
Star# AZ EL LBCR LBCR Dimm
8 382 76 90502 90457 1.25"
34 8 32 92800 92756 1.2"
17 20 52 94031 94029 1.28"
15 38 71 95300 95257 1.08"
32 42 35 100403 100400 1.25
24 51 54 101459 101456 1.57"
31 77 34 102821 102818 1.54"
25 85 59 103833 103829 1.42'
19 116 74 104732 104728 1.53"
23 146 31 110028 110025 1.08
16 175 45 111111 111109  
14 201 69 112734 112732 1.85"
5 265 50 114404 114400  
8 303 55 120159 120158 1.88
15 318 70 121234 121231  
1 338 38     Too bright, short only.

too few downloaded pixels error: 2021/09/14 11:59:59.634000 E B CAMERA CCDCTRL too few downloaded pixels (42465396 instead of 42467328) [ccdctrl.c:2761] during dofpia

12:27 Slewing for LBC skyflats.

12:31 starting g & i skyflats. Got a few at PA0 but missed the window for PA180, those are a bit on the high side.

Andrew has powered off the laser.

12:43. Thats all she wrote. Closing up.

Calibrations

Darks for transform data: 10x[1x4.0] lir luci1.20210914.0121- LUCI1 darks for transform data: 10x[1x4.0] lir luci1.20210914.0121-130 -- BarryRothberg - 13 Sep 2021Taking mods dual imaging flats and mods red only imaging flats

LBC 25 Bias.

LBC's powered off
LUCI verified safed

-- BarryRothberg - 13 Sep 2021
Topic revision: r11 - 02 Jun 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback